The perimeter of the rectangular playing field is 360 yards. The length of the field is 4 yards less than triple the width. What are the dimensions of the playing field?
Answer by GenericMIKE(1)

l = length = 3w -4
w = width
P = perimeter = 360 yd = 2(l) + 2(w)
Plug in for l in the above equation, and solve for w (width) using algebra.
360 = 2(3w - 4) + 2w
360 = 6w - 8 + 2w
368 = 8w
46 = w
So,
Using the value that we found for w, the width, we can now find the length.
l = 3w - 4
l = 3(46) - 4
l = 138 - 4
l = 134 yd